OpenStudy (anonymous):

Square root of 48x^3 divided by the square root of 3xy^2

OpenStudy (dopeeaf):

i would say 4y

OpenStudy (anonymous):

You can combine them under one square root to start. then reduce under the radical

OpenStudy (anonymous):

I thought it would be something like \[2xy \sqrt{4}\], but I'm not exactly sure..

OpenStudy (anonymous):

\[\sqrt{\frac{ 48x^3 }{ 3xy^2 }}\]

OpenStudy (anonymous):

That's what you get when you combine the square root. Reduce that to start

OpenStudy (dopeeaf):

48x^3/3xy^2 16x^2/y^2 =4x/y idk , i might be wrong

OpenStudy (anonymous):

you're right

OpenStudy (anonymous):

so its like \[\frac{ \sqrt{48x^3} }{ 3xy^2 } \] then you divide 48 by 3 to get 16, subtract x from x^3 to get x^2...

OpenStudy (dopeeaf):

you're right

OpenStudy (anonymous):

You're right, but be careful with the square root sign. Make sure it covers the whole fraction.

OpenStudy (anonymous):

okay, and @peachpi thanks for pointing that out